# Toward Honest Language Models for Deductive Reasoning

### v2

@@ 1 Introduction
Several training approaches have been widely used to enable models to perform reasoning tasks. Supervised fine-tuning (SFT) (wu2025generalization) has proven highly effective at quickly aligning models to desired behaviors, but it tends to overfit to demonstrations and struggles to generalize beyond the dataset distribution. Reinforcement learning methods such as Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) compares multiple rollouts of the same query to assign relative advantages and optimizes only for final verifiable outcomes. However, when all rollouts in the batch are incorrect and receive identical rewards, their relative advantages collapse to zero, leading to several issues including vanishing gradients and reinforcing dishonest overconfidence (liu2025understanding; yu2025dapo; zheng2025group). To address these issues, studies have explored integrating SFT with RL. However, most either lack access to a verifiable ground-truth trajectory for guidance (yan2025learning; liu2025ghpo; huang2025blending; nath2025adaptive), or do not directly address the stability of policy gradient updates, leaving them still susceptible to zero-variance issues at certain training stages (wu2025thought; chen2025step; zhang2025policy). This motivates the need for new datasets and methods that specifically target honest deductive reasoning.
@@ 3 Deductive Reasoning Dataset Construction
A first instantiation of the DAH is in the domain of linear algebra. In this case, each hyperedge reduces to a simple edge corresponding to a linear equation between two nodes. Specifically, for nodes $m,n\in V$, an edge encodes a relation of the form $am+bn=c,$ where $a,b,c\in\mathbb{Z}$. The values of root variables $r\in R$ are provided as input. If there are $k$ edges along the unique path from some root $r$ to the query node $q$, the resulting problem amounts to solving a system of $k$ linear equations to obtain the value of $q$. To make the tasks accessible to language models, we convert each equation into a natural language sentence comparing the prices of food dishes, and pose the query as a question about the price of object $q$ (ouyang2025treecut). An example of such a prompt is provided in Table 3.
@@ 5.1 Methodology: Anchor
According to Proposition 1, for every token $y_{t}^{\star}$ that is not clipped, the update direction coincides with the SFT token-level gradient $\nabla_{\theta}\log\pi_{\theta}(y_{t}^{\star}\mid x,y^{\star}_{<t})$, scaled by the positive factor $\smash{\hat{A}^{\star}w_{t}^{\star}(\theta)/(G\,|y^{\star}|)}$. Tokens for which clipping is active contribute zero. Thus, Anchor effectively augments the GRPO gradient with an SFT-like term. In the extreme case where each group contains only the ground truth rollout ($G=1$), the GRPO surrogate gradient reduces exactly to the SFT gradient on the ground-truth tokens, with clipping applied. Consequently, Anchor unifies the strengths of SFT that learns from explicit supervision of correct reasoning trajectories and GRPO explores beyond the dataset via relative credit assignment, thereby ensuring stable gradient updates even in challenging scenarios where unguided GRPO would otherwise provide no learning signal.
@@ 8 Conclusion
This work investigates honesty alignment in language models through controlled deductive reasoning tasks that do not involve human subjects or sensitive data. The datasets are generated from mathematical and logical structures, ensuring no privacy concerns, legal risks, or discriminatory content. Our methodology focuses on developing models that recognize unanswerable queries and abstain appropriately, aiming to reduce the risk of misleading outputs and improve reliability in downstream applications. While our approach seeks to mitigate harms associated with dishonest reasoning, potential misuse of more capable aligned models for deceptive purposes remains a concern. We encourage responsible research and deployment consistent with the ICLR Code of Ethics.
@@ Appendix B Limitations
Our evaluation focuses on three publicly available Qwen-based models (0.6B, 1.7B, and 3B parameters). While this choice reflects practical compute considerations, it is also deliberate: these models are widely used, span a meaningful range of capacities, and perform competitively on a broad set of recent benchmarks, making them strong, representative proxies for contemporary compact LLMs. A future extension to additional—and especially larger—architectures would be valuable to further stress-test Anchor and broaden external validity, but we expect the core trends reported here to hold given the diversity and state-of-the-art standing of the selected Qwen variants.
In addition, we focus exclusively on deductive reasoning tasks to avoid confounding effects with factual recall. This choice enables a controlled evaluation of our approach. An interesting avenue for future work is to extend the analysis to tasks that combine deductive reasoning with external factual or probabilistic knowledge, as many real-world applications demand. Such extensions would broaden the scope of our findings and provide a more comprehensive picture of the reasoning challenges faced by deployed language models.
@@ Appendix H Additional Discussion on Process Rewards
Appendix I Clarifications on LLM Usage

### v4

@@ 1 Introduction
Several training approaches have been widely used to enable models to perform reasoning tasks. Supervised fine-tuning (SFT) (wu2025generalization) has proven highly effective at quickly aligning models to desired behaviors, but it tends to overfit to demonstrations and struggles to generalize beyond the dataset distribution. Reinforcement learning methods such as Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) compares multiple rollouts of the same query to assign relative advantages and optimizes only for final verifiable outcomes. However, when all rollouts in the batch are incorrect and receive identical rewards, their relative advantages collapse to zero, leading to several issues including vanishing gradients and reinforcing dishonest overconfidence (liu2025understanding; yu2025dapo; zheng2025group). Studies have explored integrating GRPO with SFT to address these issues. However, most either lack access to a verifiable ground-truth trajectory for guidance (yan2025learning; liu2025ghpo; huang2025blending; nath2025adaptive), or do not directly address the stability of policy gradient updates, leaving them still susceptible to zero-variance issues at certain training stages (wu2025thought; chen2025step; zhang2025policy). This motivates the need for new datasets and methods that target honest deductive reasoning.
@@ 3 Deductive Reasoning Dataset Construction
A first instantiation of the DAH is in the domain of linear algebra. In this case, each hyperedge reduces to a simple edge corresponding to a linear equation between two nodes. Specifically, for nodes $m,n\in V$, an edge encodes a relation of the form $am+bn=c,$ where $a,b,c\in\mathbb{Z}$. The values of root variables $r\in R$ are provided as input. If there are $k$ edges along the unique path from some root $r$ to the query node $q$, the resulting problem amounts to solving a system of $k$ linear equations to obtain the value of $q$. To make the tasks accessible to language models, we convert each equation into a natural language sentence comparing the prices of food dishes, and pose the query as a question about the price of object $q$ (ouyang2025treecut). This natural language formulation prevents models from bypassing the intended graph traversal process by applying general linear algebra techniques, such as Cramer’s rule, directly to the mathematical form of the linear equations. An example of such a prompt is provided in Table 3.
@@ 5.1 Methodology: Anchor
According to Proposition 1, for every token $y_{t}^{\star}$ that is not clipped, the update direction $\nabla_{\theta}\log\pi_{\theta}(y_{t}^{\star}\mid x,y^{\star}_{<t})$ is scaled by the positive factor $\smash{\hat{A}^{\star}w_{t}^{\star}(\theta)/(G\,|y^{\star}|)}$, treating the ground-truth trajectory as an additional rollout. Tokens for which clipping is active contribute zero. Thus, Anchor effectively augments the GRPO gradient with an SFT-like term. In the extreme case where each group contains only the ground truth rollout ($G=1$), the GRPO surrogate gradient reduces exactly to the SFT gradient on the ground-truth tokens, with clipping applied. Consequently, Anchor unifies the strengths of SFT that learns from explicit supervision of correct reasoning trajectories and GRPO explores beyond the dataset via relative credit assignment, thereby ensuring gradient updates even in challenging scenarios where unguided GRPO would otherwise provide no learning signal.
@@ 8 Conclusion
This work investigates honesty alignment in language models through controlled deductive reasoning tasks that do not involve human subjects or sensitive data. The datasets are generated from mathematical and logical structures, ensuring no privacy concerns, legal risks, or discriminatory content. Our methodology focuses on developing models that recognize unanswerable queries and abstain appropriately, aiming to reduce the risk of misleading outputs and improve reliability in downstream applications. While our approach seeks to mitigate harms associated with dishonest reasoning, potential misuse of more capable aligned models for deceptive purposes remains a concern.
@@ Appendix B Limitations
Our evaluation focuses on three publicly available Qwen-based models. While this choice reflects practical compute considerations, it is also deliberate: these models are widely used, span a meaningful range of capacities, and perform competitively on a broad set of recent benchmarks, making them strong, representative proxies for contemporary compact LLMs. Extending evaluations to substantially larger models would not offer a commensurate scientific benefit considering the cost. A future extension to additional architectures would be valuable to further stress-test Anchor and broaden external validity, but we expect the core trends reported here to hold given the diversity and state-of-the-art standing of the selected Qwen variants.
In addition, we focus exclusively on deductive reasoning tasks to avoid confounding effects with factual recall. Our goal in this paper is not to improve performance on compositional realistic dataset benchmarks, but to understand and study one form of honesty: whether a model can recognize when premises are insufficient for a conclusion in multi-step deductive reasoning. This choice offers the crucial advantage of controllability: we can define tasks where the answerability is fully determined by the underlying graph structure, precisely manipulate difficulty and reasoning depth, and ensure clean separation between reasoning and knowledge. Realistic datasets tend to mix multiple skills: reasoning, domain knowledge, linguistic expectations, and even stylistic cues, making it unclear what type of error a model makes when it fails. They could also introduce contamination risks because pretraining corpora may include similar examples. However, an interesting but long-term avenue for future work is to extend the analysis to tasks that combine deductive reasoning with external factual or probabilistic knowledge, as many real-world applications demand. Such extensions would provide a more comprehensive picture of the reasoning challenges faced by deployed language models.
The zero-variance issue we address arises only in policy-gradient algorithms that use group-relative advantages, such as GRPO, and does not apply to prior RL approaches like PPO (schulman2017proximal).
@@ Appendix H Additional Discussion on Process Rewards
Appendix I Summary of Key Takeaways
Our contributions extend beyond a single methodological insight. First, we design controlled, multi-step deductive reasoning datasets specifically tailored for studying honest reasoning behavior. We analyze why existing approaches such as SFT, GRPO, and various patching strategies on them fail in this setting. We provide theoretical analysis showing how Anchor, injecting ground-truth reasoning trajectories, stabilizes policy gradient updates by mitigating gradient vanishing, and we demonstrate empirically that this approach leads to substantial improvements. To our knowledge, no prior work addresses all of these components together. The simplicity of our method is intentional, not a drawback: if a simple mechanism can resolve instability and enable models to learn long-range deductive patterns, introducing additional modules or constraints would add unnecessary complexity and overfit the datasets tested.
In summary, our paper introduces the concept of honesty in deductive reasoning, constructs two controllable multi-step deductive reasoning datasets, analyzes the limitations of existing approaches such as SFT and GRPO, and proposes Anchor: a simple yet effective method for unifying SFT and GRPO signals by injecting ground-truth trajectories into policy rollouts. Through theoretical insights and extensive empirical evidence, we show that Anchor stabilizes training, mitigates gradient vanishing, and significantly improves both deductive accuracy and honest abstention.
Appendix J Clarifications on LLM Usage